Use the frequency distribution shown below to construct an High Temperatures (F) Class expanded frequency distribution 51-61 62-72 73-83 18-28 29-39 40-50 84-94 Frequency, f 18 45 68 69 75 68 22 Complete the table below. High Temperatures (°F) Class Frequency, f Midpoint frequency (Round to the nearest hundredth as needed.) Cumulative frequency Relative 18-28 5.2 18 18
Use the ogive to approximate (a) the number in the sample, (b) the location of the greatest increase in frequency (a) The approximate number in the sample is Male Beagles 50- 45- 40- 35- 30- 25- 20- 15- 10- 5- 0 18.5 21.5 24.5 27.5 30.5 33.5 Weight (in pounds) Aouanbau angeinuuro 오오호호호우 후호
Answer #1

1)

The relative frequencies are obtained using the formula,

frequency Relative frequency Total frequency

The cumulative frequencies are obtained by adding current class frequency with the sum of its predecessors

Class Frequency Midpoint Relative frequency Cumulative frequency
18-28 18 23 18/365=0.049 0.049
29-39 45 34 45/365=0.123 0.049+0.123=0.173
40-50 68 45 68/365=0.186 0.173+0.186=0.359
51-61 69 56 69/365=0.189 0.359+0.189=0.548
62-72 75 67 75/365=0.205 0.548+0.205=0.753
73-83 68 78 68/365=0.186 0.753+0.186=0.940
84-94 22 89 22/365=0.060 0.940+0.060=1
SUM 365
Class Frequency Midpoint Relative frequency Cumulative frequency
18-28 18 23 0.049 0.049
29-39 45 34 0.123 0.173
40-50 68 45 0.186 0.359
51-61 69 56 0.189 0.548
62-72 75 67 0.205 0.753
73-83 68 78 0.186 0.940
84-94 22 89 0.060 1
SUM 365

2)

The frequencies are obtained by counting the number of points falls in the class.

Class Frequency Relative frequency Cumulative frequency
18.5-21.5 4 0.25 0.25
21.6-24.5 3 0.1875 0.4375
24.6-27.5 3 0.1875 0.625
27.6-30.5 3 0.1875 0.8125
30.6-33.5 3 0.1875 1
SUM 16

a)

Total number in the sample = 16

b)

Greatest increase in frequencies is in class 18.5-21.5